# “Stones from Other Hills can Polish Jade”: Zero-shot Anomaly Image Synthesis via Cross-domain Anomaly Injection

### v1

@@ 4.1 Experimental Setup
CutPaste 93.9 47.8 79.9 93.6 22.6 76.6 92.7 39.5 80.6
NSA 93.2 50.2 82.7 91.8 19.1 75.8 94.9 36.7 87.1
DRAEM 92.0 54.6 83.4 91.6 26.9 79.1 94.1 43.8 85.6
RealNet 84.2 42.1 73.2 90.9 21.0 80.7 81.5 22.6 68.8
GLASS 92.9 56.8 84.4 92.4 24.5 78.7 94.9 47.1 87.9
CAI 96.1 62.3 89.7 93.6 29.9 83.1 98.2 70.0 94.4
Evaluation Protocol. To conduct a dedicated evaluation on ZSAS, our evaluation protocol is designed based on [hu2024anomalydiffusion], the latest work that has conducted a specialized evaluation of anomaly synthesis. Concretely, pseudo anomalies are evaluated in terms of two facets: (1) Utility in IAD. Since the goal of ZSAS is to serve IAD here, we evaluate the utility of different ZSAS solutions’ pseudo anomalies by a unified IAD framework: With each target image from a normal training set, ten pseudo anomaly images are synthesized. Then, a binary-class training set is built by combining given normal images and pseudo anomalies. Since our goal is to manifest the performance difference of various ZSAS solutions rather than pursuing SOTA performance on a certain dataset, we only train a simple UNet from [zavrtanik2021draem, hu2024anomalydiffusion] by Focal loss [ross2017focal], and avoid introducing more modules or boosting tricks. Following [hu2024anomalydiffusion], with the pixel-level anomaly mask output by the trained UNet, IAD performance is evaluated by three metrics: Pixel-level Area under Receiver Operating Characteristic Curve (AUC), Average Precision (AP) and Per-Region Overlap (PRO). For each ZSAS solution, we select the best-performing model among 200 training epochs for comparison to avoid setting optimal epoch number. The training process of all ZSAS solutions shares the same hyperparameters and random seed (see supplementary material). (2) Authenticity. We measure authenticity by FID [heusel2017gans], which reflects the similarity of pseudo anomalies and domain-specific anomalies in the test set.
Benchmark Datasets and Competing Methods. We use three widely-used IAD benchmark datasets: MvTec-AD [bergmann2019mvtec], VisA [zou2022spot] and KSDD2 [bovzivc2021mixed]. Since MvTec-AD and VisA are included into our domain-agnostic anomaly dataset, it should be noted that they will be excluded from ZSAS when experiments are conducted on themselves to avoid information leakage. As to competing ZSAS methods, we compare the proposed CAI with 5 SOTA ZSAS solutions: CutPaste [li2021cutpaste], NSA [schluter2022natural], DRAEM [zavrtanik2021draem], GLASS [chen2024unified] and RealNet [zhang2024realnet]. Besides, we also include the latest FSAS solution AnomalyDiffusion (ADiff) [hu2024anomalydiffusion] as a reference. More details of experimental setup are given in supplementary material.
4.2 Experimental Results
Figure 5: Pseudo anomalies of CAI (row 1) and CDM (row 2-5).
@@ 4.2 Experimental Results
CutPaste NSA DRAEM RealNet GLASS CAI
MvTec-AD 123.3 97.7 164.7 143.4 168.3 102.2
VisA 61.7 34.5 87.7 86.9 88.7 31.2
KSDD2 151.3 166.3 252.9 301.2 221.2 136.9
Quality of Pseudo Anomalies. We first compare the quality of pseudo anomalies from different ZSAS solutions in Table 1-2. For utility in IAD, pseudo anomalies synthesized by CAI constantly achieve the best IAD performance in terms of all metrics on all datasets (see Table 1). Compared with SOTA ZSAS counterparts, CAI brings notable improvement in almost all cases, while it even outperforms the 2nd best performer by 22.9% AP and 6.5% PRO on KSDD2. Such results verify the effectiveness of CAI for ZSAS in IAD. As to authenticity, pseudo anomalies synthesized by CAI attain the lowest FID on VisA and KSDD2 dataset, and the second lowest FID on MvTec-AD (see Table 2), despite that NSA and CutPaste actually enjoy an unfair advantage in FID calculation: They sample normal image patches of the train set as anomaly patterns, which are naturally similar to the test set of the same domain. By contrast, without using patches of the same domain as anomalies, CAI’s FID is comparable or superior to NSA and CutPaste, which reveals that CAI can synthesize highly authentic pseudo anomalies for the specific domain. In addition, we also have two interesting observations: First, CutPaste’s FID is higher than NSA, which can be ascribed to image artifacts caused by pasting-based anomaly injection. Second, although model-based RealNet yields lower FID than model-free methods DRAEM/GLASS on MvTec-AD/VisA dataset, it performs poorly on KSDD2 dataset, whose images have evidently different characteristics from MvTec-AD/VisA. Such results also reveal the complexity in designing the tricky generative process for various domains. As a summary, CAI can synthesize high-quality pseudo anomalies in a simple and training-free manner.
@@
(d) Different ratios of scale.
@@
Ablation Studies. In Table 3, we conduct ablation studies to verify the role of each technique in CAI: (1) Without PE-based injection (line 1), the performance is degraded by over $5\%$ under all metrics, which unveils its importance in eliminating artifacts. (2) Using MAS constantly outperforms the cases where only one scale is used (line 2-4), which necessitates synthesizing multi-scale pseudo anomalies. (3) Despite of fairly close AUC and AP, STM’s removal leads to more tangible loss in PRO (line 5). The loss will be more severe when small anomalies take a larger proportion in the anomaly dataset. (4) Injection location selection (ILS) has a crucial impact on performance, the absence of which causes over 10% loss under each metric (line 6).
Table 3: Results (%) of ablation studies for CAI.
PE Large Medium Small STM ILS AUC AP PRO
- 90.5 50.9 81.8
- - 96.0 61.5 88.6
- - 95.4 60.4 88.2
- - 95.0 60.4 88.2
- 96.2 62.4 89.3
- 85.2 42.3 74.9
96.1 62.3 89.7
Discussion. (1) FSAS vs. ZSAS: We also compare ZSAS solutions (CAI and GLASS [chen2024unified]) with the recent FSAS solution ADiff [hu2024anomalydiffusion]. ADiff is a strong FSAS competitor, as it exploits domain-specific anomalies from all anomaly classes of the test set as training data. As shown in Fig. 6(b), although the performance gap still exists, CAI has evidently narrowed the gap between ZSAS and FSAS when compared with the latest GLASS. (2) Data sources. We collect cross-domain anomalies from three sources: Online search (OS), reality photographing (RP) and existing datasets (ED), while images from OS and RP are newly collected. We compare three data source configurations: OS+RP+ED, OS+RP and ED only. As shown in Fig. 6(c), three configurations actually achieve comparable performance, which suggests that real anomalies from different sources can be readily transferred to a specific domain, while this also verifies the quality of our newly-collected anomaly data. (3) Ratio of large/medium/small anomalies. In addition to the 4:3:3 ratio that we use by default, we also test the ratio of 6:2:2, 2:6:2 and 2:2:6, which place more emphasis on large/medium/small pseudo anomalies respectively. As shown by Fig. 6(d), results suggest that the performance is actually insensitive to ratio change. Due to page limit, more results are given in supplementary material.

### v2

@@ 4.1 Experimental Setup
Evaluation Protocol. To conduct a dedicated evaluation on ZSAS, our evaluation protocol is designed based on [hu2024anomalydiffusion], the latest work that has conducted a specialized evaluation of anomaly synthesis. Concretely, pseudo anomalies are evaluated in terms of two facets: (1) Utility in IAD. Since the goal of ZSAS is to serve IAD here, we evaluate the utility of different ZSAS solutions’ pseudo anomalies by a unified IAD framework: With each target image from a normal training set, ten pseudo anomaly images are synthesized. Then, a binary-class training set is built by combining given normal images and pseudo anomalies. Since our goal is to manifest the performance difference of various ZSAS solutions rather than pursuing SOTA performance on a certain dataset, we only train a basic UNet from [zavrtanik2021draem, hu2024anomalydiffusion] by Focal loss [ross2017focal], and avoid introducing more modules or boosting tricks. Following [hu2024anomalydiffusion], with the pixel-level anomaly mask output by the trained UNet, IAD performance is evaluated by three metrics: Pixel-level Area under Receiver Operating Characteristic Curve (AUC), Average Precision (AP) and Per-Region Overlap (PRO). For each ZSAS solution, we select the best-performing model among 200 training epochs for comparison to avoid setting optimal epoch number. The training process of all ZSAS solutions shares the same hyperparameters and random seed (see supplementary material). (2) Authenticity. We measure authenticity by FID [heusel2017gans]: Lower FID reflects better similarity between pseudo anomalies and domain-specific anomalies of the test set.
Benchmark Datasets and Competing Methods. We use three widely-used IAD benchmark datasets: MvTec-AD [bergmann2019mvtec], VisA [zou2022spot] and KSDD2 [bovzivc2021mixed]. Since MvTec-AD and VisA are included into our domain-agnostic anomaly dataset, it should be noted that they will be excluded from ZSAS when experiments are conducted on themselves to avoid information leakage. As to competing ZSAS methods, we compare the proposed CAI with 5 SOTA ZSAS solutions: CutPaste [li2021cutpaste], NSA [schluter2022natural], DRAEM [zavrtanik2021draem], GLASS [chen2024unified] and RealNet [zhang2024realnet]. Since NSA also employs Poisson editing, it is also categorized into NSA-N and NSA-M, like our CAI-N and CAI-M. Besides, we also include the latest FSAS solution AnomalyDiffusion (ADiff) [hu2024anomalydiffusion] as a reference. More details of experimental setup are given in supplementary material.
4.2 Experimental Results
Figure 5: Pseudo anomalies of CAI (row 1) and CDM (row 2-5).
@@ 4.2 Experimental Results
CutPaste 89.6 38.5 71.9 93.0 15.8 72.2 89.0 39.6 70.9
NSA-N 92.4 49.1 80.6 95.7 37.3 83.9 80.5 4.5 55.5
NSA-M 91.2 47.2 78.8 94.5 25.9 83.4 87.6 34.8 69.9
DRAEM 91.0 54.5 82.5 93.2 37.3 85.6 93.8 52.9 86.6
RealNet 85.1 39.0 70.9 85.5 21.8 79.1 81.3 18.9 64.3
GLASS 91.1 52.0 82.3 91.4 27.0 79.1 92.5 42.7 82.3
CAI-N 95.6 62.6 87.5 96.1 39.7 87.7 96.6 70.4 89.6
CAI-M 94.7 61.6 86.6 95.9 40.1 89.1 96.8 66.3 89.6
Quality of Pseudo Anomalies. We first compare the quality of pseudo anomalies from different ZSAS solutions in Table 1-2. For utility in IAD, pseudo anomalies synthesized by CAI constantly achieve the best IAD performance in terms of all metrics on all datasets (see Table 1). Meanwhile, CAI-N (Normal mode PE) achieves better overall performance than CAI-M (Mixed mode PE). Compared with SOTA ZSAS counterparts, CAI brings notable improvement in almost all cases, while it even outperforms the 2nd best performer by 17.5% AP on KSDD2. Such results verify the effectiveness of CAI for ZSAS in IAD. As to authenticity, pseudo anomalies synthesized by CAI-M and CAI-N attain the lowest and 2nd lowest FID on all 3 datasets (see Table 2), even though NSA and CutPaste actually enjoy an unfair advantage in FID calculation: They sample normal image patches of the train set as anomaly patterns, which are naturally similar to the test set of the same domain. By contrast, without using patches of the same domain as anomalies, CAI’s FID is superior to NSA and CutPaste, which reveals that CAI can synthesize highly authentic pseudo anomalies for the specific domain. In addition, we also have two interesting observations: First, CutPaste’s FID is higher than NSA, which can be ascribed to image artifacts caused by pasting-based anomaly injection. Second, although model-based RealNet yields lower FID than model-free methods DRAEM/GLASS on MvTec-AD/VisA dataset, it performs poorly on KSDD2 dataset, whose images have evidently different characteristics from MvTec-AD/VisA. Such results also reveal the complexity in designing the tricky generative process for various domains. As a summary, CAI can synthesize high-quality pseudo anomalies in a simple and training-free manner.
Figure 6: IAD detector trained by pseudo anomalies synthesized by CAI and DAAD can generalize to different unseen anomalies, e.g. wrong pill type, bent grid, transistor misplacement.
@@
CutPaste NSA-N NSA-M DRAEM RealNet GLASS CAI-N CAI-M
MvTec-AD 153.4 138.4 136.9 182.6 139.5 175.6 121.8 107.0
VisA 98.5 78.1 74.9 99.3 81.7 93.8 45.9 37.9
KSDD2 201.9 187.0 185.2 244.9 301.2 224.9 169.4 147.9
Generalization to Unseen Anomalies. As discussed above, the anomaly categories in DAAD dataset are still finite. Thus, one may wonder whether pseudo anomalies synthesized by CAI and DAAD dataset can generalize to unseen anomaly categories. For example, on MvTec-AD dataset, bent wire and accessory misplacement are not included in our DAAD dataset. However, as shown by Fig. 6, the IAD detector, which is trained by pseudo anomalies synthesized by our CAI and DAAD datasets, can effectively detect those unseen anomalies. This demonstrates that CAI enables the detector to generalize to unseen anomalies.
@@
(d) Different sampling number.
@@
Ablation Studies. In Table 3, we conduct ablation studies to verify the role of each technique in CAI-N: (1) Without PE-based injection (line 1), the performance is degraded by over $5\%$ under all metrics, which unveils its importance in eliminating artifacts. (2) The removal of Source-target matching (STM) leads to tangible performance loss in terms of all three metrics (line 2), which can be up to 0.8%. This verifies the necessity of STM. (3) Injection location selection (ILS) provides a more precise and valid injection position, enhancing performance by up to 1.6% (line 3).
Discussion. (1) We first compare the cases where different ratios of large, medium and small pseudo anomalies are used for CAI: Large-biased (6:2:2), medium-biased (2:6:2), small-biased (2:2:6) and balanced (4:3:3). As shown in Table 4, the balanced case achieves the better overall performance than any of the biased cases. This manifests the importance to perform a muti-scale anomaly synthesis that considers all scales fairly. (2) FSAS vs. ZSAS: We also compare ZSAS solutions (CAI and GLASS [chen2024unified]) with the recent FSAS solution ADiff [hu2024anomalydiffusion]. ADiff is a strong FSAS competitor, as it exploits domain-specific anomalies from all anomaly classes of the test set as training data. As shown in Fig. 7(b), although the performance gap still exists, CAI has evidently narrowed the gap between ZSAS and FSAS when compared with the latest GLASS. (3) Data sources. We collect cross-domain anomalies from three sources: Online search (OS), reality photographing (RP) and existing datasets (ED), while images from OS and RP are newly collected. We compare three data source configurations: OS+RP+ED, OS+RP and ED only. As shown in Fig. 7(c), three configurations actually achieve comparable performance, which suggests that real anomalies from different sources can be readily transferred to a specific domain, while this also verifies the quality of our newly-collected anomaly data. (4) Number of UNet detector sampling. We adjusted the number of samples used in training the UNet detector, reporting scenarios with 500 and 1500 samples in CAI. As shown in Fig. 7(d), it was observed that the model’s performance does not undergo drastic change, so we simply choose 1000 by default. (5) Computational cost. It takes an average 0.019s and 0.018s for CAI-N and CAI-M to synthesize one pseudo anomaly image on the MvTec-AD dataset, which are fairly acceptable. Due to page limit, more results on cost are given in supplementary material.
Table 3: Results of ablation studies on core components of CAI.
PE STM ILS AUC AP PRO
- 89.7 50.1 77.5
- 95.0 61.8 86.7
- 95.1 61.3 85.9
95.6 62.6 87.5
Table 4: Discussion on multi-scale anomaly synthesis.
AUC AP PRO
Large-biased (6:2:2) 95.2 61.7 86.8
Medium-biased (2:6:2) 94.8 61.5 86.7
Small-biased (2:2:6) 95.4 61.8 87.8
Balanced (4:4:3) 95.6 62.6 87.5